# Unharmful Backdoor-based Client-side Watermarking in Federated Learning

### v1

@@ Abstract
Protecting intellectual property (IP) in federated learning (FL) is increasingly important as clients contribute proprietary data to collaboratively train models. Model watermarking, particularly through backdoor-based methods, has emerged as a popular approach for verifying ownership and contributions in deep neural networks trained via FL. By manipulating their datasets, clients can embed a secret pattern, resulting in non-intuitive predictions that serve as proof of participation, useful for claiming incentives or IP co-ownership. However, this technique faces practical challenges: client watermarks can collide, leading to ambiguous ownership claims, and malicious clients may exploit watermarks to inject harmful backdoors, jeopardizing model integrity. To address these issues, we propose Sanitizer, a server-side method that ensures client-embedded backdoors cannot be triggered on natural queries in harmful ways. It identifies subnets within client-submitted models, extracts backdoors throughout the FL process, and confines them to harmless, client-specific input subspaces. This approach not only enhances Sanitizer’s efficiency but also resolves conflicts when clients use similar triggers with different target labels. Our empirical results demonstrate that Sanitizer achieves near-perfect success in verifying client contributions while mitigating the risks of malicious watermark use. Additionally, it reduces GPU memory consumption by 85% and cuts processing time by at least 5$\times$ compared to the baseline. Our code is open-sourced at https://anonymous.4open.science/r/Sanitizer-5933.
@@ 1 Introduction
Federated Learning (FL) [1] is on its way to becoming a standard for training deep learning models with data distributed across clients. As the training data and the resultant model are of high commercial value [2, 3, 4, 5, 6, 7], protecting their intellectual property (IP) has become particularly important [8]. To this end, backdoor attacks can serve as a positive means to embed a unique watermark into the model for ownership verification. During model training, one could train the model to recognize a special pattern known as a trigger, such that when the trigger is attached to the input, the model returns a predefined output regardless of the actual content [9, 10]. The special pattern is often kept secret; hence, the demonstration of such non-intuitive behavior can be used as a watermark to claim the ownership or contributions.
Figure 1: Without Sanitizer, a malicious client (e,g., Bob) can control the model by predicting a stop sign attached with a special trigger (originally used as a watermark) to be the “Ahead Only”. Watermark collision may also occur when two clients (e,g., Sam and Jack) use a similar trigger but designate different target labels.
@@
  • Potential Misuse for Malicious Purposes: Malicious clients can exploit watermarks to control model predictions in harmful ways [16, 17, 18], such as predicting a stop sign attached with a special pattern originally used as a watermark to be “Ahead Only.”
@@
  • Watermark Collisions: As clients (benign or malicious) choose their own trigger-output pair as the watermark, a collision may occur when two clients use a similar pattern as the trigger but designate different outputs (i.e., class labels). Such a collision can lead to both watermarks being unlearnable.
@@
Overall, the use of backdoor attacks for client-side watermarking not only poses a serious security threat to FL but may also become an obstacle to IP protection in FL.
Figure 2: Sanitizer makes the trigger ineffective when placed on a natural image (e.g., Bob). But it enables trigger to output the target label as a watermark for IP protection only when placed in the client-specific unharmful environment (e.g., Sam and Jack), and no collision even if they use a similar trigger with different target labels.
@@
To alleviate the above problems, we explore a server-side sanitization process called Sanitizer, with an overarching idea of making all the implanted backdoors, whether they are benign or malicious, effective only in an unharmful environment. Sanitizer defines a client-dependent input subspace composed of unnatural samples and moves the backdoors implanted by each client to its designated subspace. As a server-side method, clients do not need to actively participate in Sanitizer and can embed watermarks as usual. As shown in Figure 2, Sanitizer makes the trigger ineffective when placed on a normal image, but it activates the backdoor and outputs a predefined target class when placed in the client-specific unharmful environment (i.e., a unique-color unnatural picture for this example) while mitigating potential backdoor conflicts among clients.
@@
Challenges. A straightforward approach involves reverse engineering the hidden triggers embedded inside each client-submitted model at each round, unlearning them, and re-establishing the trigger-output associations in an unharmful environment post-FL. However, referred to as the baseline method in this paper, it is inefficient as reverse engineering at each round incurs significant resource and time consumption, making it impractical for FL applications, as shown in Figure 3 (red). To optimize the efficiency for FL, we need to minimize server-side processing time in each round to prevent delay in subsequent rounds during FL training, and we also need to ensure a low resource consumption to enable parallel trigger recovery across multiple client-submitted models.
@@
Figure 3: Sanitizer offers a significantly better scalability. It keeps the total server-side time consumption to run a 10-round FL on CIFAR10 with ResNet18 consistently lower (green) than the baseline (red) as the number of participating clients increases.
@@
In this paper, Sanitizer aims to efficiently sanitize backdoors for unharmful client-side watermarking. To tackle the challenges mentioned above, we propose to first extract a small backdoor subnet from each client-submitted model for reverse engineering, driven by the observation that backdoors are often encapsulated in a subset of neurons. Then, we take advantage of the iterative learning paradigm in FL and propose a lightweight method to recover the trigger implanted in each client-submitted model by spreading the reverse optimization across communication rounds. The trigger and target class gradually emerge and take shape over time. Based on the above two improvements, Sanitizer is efficient, requires minimal GPU resources, and does not hamper the overall FL training process [19]. As shown in Figure 3, Sanitizer maintains consistently low time consumption (green). With only 25 clients, the baseline (red) takes approximately five times longer than Sanitizer. Given the fixed GPU memory of the server, the disparity between the two methods becomes more pronounced as the number of clients increases. Due to lower memory consumption, Sanitizer enables the server to process more client-submitted models in parallel. After completing the FL training, we employ an unharmful relearning process to “unbind” the trigger from the clean images and re-establish its association with the unique, unharmful environment of each client. Since each client-submitted model is processed independently before aggregation at each round and each client’s trigger is ultimately confined to its own unharmful client-specific input subspace, watermark conflicts will not arise during the verification phase. In summary, our key contributions are as follows:
@@
  • We introduce a server-side sanitization process that mitigates potential harmful use in backdoor-based client-side watermarking by enabling the triggering of counter-intuitive model behavior in unharmful environments.
@@ 2 Related Work
With AI Trust emerging as a critical industry concern, as noted in Gartner’s latest technology report [20], protecting IP in both centralized and FL-trained model remains a significant ongoing research focus driven by the increasing commercial and legal demands [21, 22, 8].
@@
Watermarking in Centralized Model. Many watermarking approaches have been proposed to protect the IP of DNN model. Broadly speaking, these methods can be categorized into two classes: the feature-based [23, 24, 25, 26, 27], requiring manipulating the loss function and model parameters, and the backdoor-based [28, 29, 30, 31, 32], involving injecting a trigger set with predefined output (known only to the model owner) as the watermark. Basically, these strategies for centralized DNN models lay the groundwork for the development of watermarking in FL, where the process is more complex due to the client-server learning protocol and specific privacy requirements [33, 34, 13].
@@
Watermarking in FL-trained Model. Watermarking methods in FL can be generally categorized into server-side [34, 17] and client-side schemes, owing to the collaborative nature of client-server workflow [13]. We emphatically investigate the current client-side solutions of watermarking for IP protection in FL. One of the representative methods is FedIPR [9], which permits each client to embed their own backdoor-based and feature-based watermark into the model for contribution demonstration without exposing either their private watermark or training data to other clients. FedSOV [35] presents a cryptographic signature-based approach on top of FedIPR, allowing numerous clients to verify the ownership credentials through unforgeable digital signatures. FedCIP [33] is a feature-based watermarking framework that allows for traitor tracking while maintaining compatibility with FL security aggregation such as parameter encryption [36]. Merkle-Sign [37] are also used for client-side watermarking, which designs a security mechanism for distributed storage to protect both privacy and ownership. FedZKP [38] proposes a verification protocol based on the zero-knowledge proof for secure model ownership verification. Besides, [15] and [11] introduce a backdoor-based watermarking scheme for IP protection within a homomorphically encrypted FL framework using homomorphic encryption [36, 39] while preventing the watermarking updates from leakage.
@@
Research Gap. However, we argue that existing client-side approaches have certain limitations that warrant consideration to ensure comprehensive IP protection in FL. First, malicious clients can embed harmful misclassification rule as a malicious backdoor into the model under the guise of backdoor-based [16] watermarking, thereby compromising the model integrity and posing a significant security threat [17, 18] to real-world applications. Second, similar backdoor trigger pattern with different target labels of any two clients is likely to cause collisions, resulting in both watermarks becoming unlearnable [14, 40]. In contrast, to address these limitations, Sanitizer provides a client-agnostic server-side sanitization pipeline, not only makes the client-side watermarks unharmful but also resolves conflicts when different clients use a similar trigger with different target labels as the watermark.
@@ 3 Preliminaries
Backdoors can evidently be utilized for both benign and malicious purposes. In this paper, when distinguishing between different purposes, we use the term “watermark” for backdoors serving benign purposes such as IP protection, and the term “harmful backdoor” or “malicious backdoor” for those serving malicious ones such as model control. Fundamentally, both purposes represent the same technique—a specific trigger-output pair embedded into a model. We use the term “backdoor” when uniformly referring to the backdoor itself without specifying purposes. These terms are used consistently throughout the paper.
In this section, we give the threat model (Section 3.1), clarify the capabilities (Section 3.2) and goals (Section 3.3) of Sanitizer, and define the poisoning formulation and unharmful setting in our context (Section 3.4).
@@ 3.1 Threat Model
We consider the threat model, where adversaries from malicious clients may launch the harmful backdoors instead of watermarks during the watermark embedding phase. In our setting, clients are authorized to embed legitimate backdoor-based watermarks into models as proof of co-ownership or contribution during the local training phase. However, adversaries on malicious clients may exploit this privilege to unlawfully implant harmful backdoor rather than watermark, and finally, control the model during inference phase for harmful purposes. It is straightforward and accessible to every client, simply placing a trigger pattern on normal data and designating a harmful target label, in line with  [16, 28, 41]. Thus, the security of backdoor-based client-side watermarking is not guaranteed. Additionally, all clients do not cooperate with the server-side sanitization process.
@@ 3.2 Capabilities of Sanitizer
Consistent with prior works [34, 17, 42, 1], the server is considered a trusted and honest party employed by the collaborating clients to facilitate FL. Sanitizer runs on the server-side and has access to the compromised models received from clients. However, it cannot interfere with the standard local training process, modify client-side operations, or access the full local dataset of individual clients. Moreover, Sanitizer has no prior knowledge about the potential backdoor triggers or target classes, and does not distinguish between benign and harmful backdoors. Similar to most existing defenses [43, 44, 45, 46, 47, 48], Sanitizer can only get access to a limited small portion of reserved clean dataset as defense data, which is common and only designed to drive the sanitization mechanisms. We exclude the defense data from all clients’ local training dataset, demonstrating that there is no overlap between them.
[…]
Figure 5: Examples of harmful (natural) inputs compared to the unharmful (artificial) inputs of each client (e.g., Bob, Sam, and Jack).
@@ 3.3 Goals of Sanitizer
  • Effectiveness. Effectiveness implies that Sanitizer makes each client deterministically demonstrate their co-ownership or contribution to the FL-trained model by achieving near-perfect success in verifying the watermark in their respective unharmful environment, while ensuring no conflicts arise among them.
@@
  • Fidelity. Fidelity refers to the requirement that Sanitizer should have negligible impact on the main functionality and integrity of the FL-trained model, maintaining high accuracy on clean test samples.
  • Efficiency. The sanitization process should run in an efficient manner during the FL process, characterized by the low time and resource consumption on the server-side, two crucial factors in this context.
@@ 3.4 Notations
in which $\mathcal{D}_{c}$ is the clean dataset of client $k$, and $\mathcal{D}_{t}$ is its trigger set for benign or malicious purposes. Besides, $f$ denotes the standard classifier parameterized by $\theta_{k}$. $\mathcal{L}$ denotes a standard loss function, e.g., cross-entropy.
Unharmful Setting. We consider the unharmful dataset $\mathcal{D}_{u}$ as a collection of unnatural images that lack practical significance and have no impact on the main task of the FL-trained model. Note that $\mathcal{X}_{u}$ must satisfy $\mathcal{X}_{u}\cap\mathcal{X}_{c}=\emptyset$, where $\mathcal{X}_{u}$ denotes the artificial input subspace and $\mathcal{X}_{c}$ denotes the clean input space, indicating no overlap between them, and ensuring $\mathcal{X}_{u}$ is distinct and significantly distant from the main task. In our setting, all the watermarks and backdoors should be transformed into an unharmful environment (i.e., being treated equally), which is artificially and uniquely tailored for each client. For example, as shown in Figure 5, we use a set of distinct and unnatural inputs in unique colors as the unharmful environments, compared to the harmful (natural) inputs.
@@ 4.1 Overview
Outline. Sanitizer is accomplished by three procedures during FL process: first, we identify and extract the backdoor subnet (Section 4.2 and Section 4.3), followed by an efficient inversion of the backdoor trigger (Section 4.4), and simultaneously perform pruning and aggregation (Section 4.5) after the identification of backdoor subnet. After completing the FL process, we conduct a final step of unharmful relearning (Section 4.6) to achieve the final unharmful effectiveness of Sanitizer. We present the details in the following sections, exemplified using a specific client $k$ at a certain round $r+1$; the approach remains consistent and independent across all clients.

### v2

# Harmless Backdoor-based Client-side Watermarking in Federated Learning

@@ Abstract
Protecting intellectual property (IP) in federated learning (FL) is increasingly important as clients contribute proprietary data to collaboratively train models. Model watermarking, particularly through backdoor-based methods, has emerged as a popular approach for verifying ownership and contributions in deep neural networks trained via FL. By manipulating their datasets, clients can embed a secret pattern, resulting in non-intuitive predictions that serve as proof of participation, useful for claiming incentives or IP co-ownership. However, this technique faces practical challenges: (i) client watermarks can collide, leading to ambiguous ownership claims, and (ii) malicious clients may exploit watermarks to manipulate model predictions for harmful purposes. To address these issues, we propose Sanitizer, a server-side method that ensures client-embedded backdoors can only be activated in harmless environments but not natural queries. It identifies subnets within client-submitted models, extracts backdoors throughout the FL process, and confines them to harmless, client-specific input subspaces. This approach not only enhances Sanitizer’s efficiency but also resolves conflicts when clients use similar triggers with different target labels. Our empirical results demonstrate that Sanitizer achieves near-perfect success verifying client contributions while mitigating the risks of malicious watermark use. Additionally, it reduces GPU memory consumption by 85% and cuts processing time by at least 5$\times$ compared to the baseline. Our code is open-sourced at https://hku-tasr.github.io/Sanitizer/.
@@ 1 Introduction
Federated Learning (FL) [1] is on its way to becoming a standard for training deep learning models with data distributed across clients. As the training data and the resultant model are of high commercial value [2, 3, 4, 5, 6, 7], protecting their intellectual property (IP) has become particularly important [8, 9]. To this end, backdoor attacks can serve as a positive means to embed a watermark into the model for ownership verification. During model training, one could train the model to recognize a special pattern known as a trigger, such that when the trigger is attached to the input, the model returns a predefined output regardless of the actual content [10, 11, 12, 13]. The special pattern is often kept secret; hence, the demonstration of such non-intuitive behavior can be used as a watermark to claim ownership or contributions.
Figure 1: Without Sanitizer, watermark collision may occur when two clients (e.g., Jack and Sam) use a similar trigger with different target labels. Furthermore, an adversary (e.g., Bob) can control the model to mispredict a query attached with a special trigger, originally used as a watermark, for malicious purposes (e.g., “Stop” becomes “Ahead Only”).
Figure 2: With Sanizier, triggers become ineffective when placed on natural images (e.g., Bob). They do not suffer from watermark collision and can only lead to non-intuitive predictions when used in client-specific harmless environments (e.g., Jack and Sam).
@@
  • Watermark Collisions: As clients choose their own trigger-output pair as the watermark, a collision may occur when two clients use a similar pattern as the trigger but designate different outputs (i.e., class labels). Such a collision can lead to both watermarks being unlearnable.
  • Potential Abuse for Malicious Purposes: Malicious clients can exploit watermarks to control model predictions in harmful ways [19, 20, 21, 22, 23], such as predicting a stop sign attached with a special pattern originally used as a watermark to be “Ahead Only.”
@@
Overall, the use of backdoor attacks for client-side watermarking can not only lead to failed or ambiguous ownership claims but also pose a serious security threat.
@@
To alleviate the above problems, we explore a server-side sanitization process called Sanitizer, with an overarching idea of making all the implanted backdoors, whether benign or malicious, effective only in a harmless environment. Sanitizer designs a client-specific input subspace composed of unnatural queries and moves the backdoors implanted by each client to its designated subspace. As a server-side method, clients do not need to actively participate in Sanitizer and can embed watermarks as usual. As shown in Figure 2, Sanitizer makes the trigger ineffective when placed on a normal image. It can only activate the backdoor and lead to a predefined target prediction when placed in the client-specific harmless environment (for this example, an unnatural image with the client’s name and a solid color background). Such a design mitigates potential backdoor conflicts among clients.
@@
Challenges. A straightforward approach involves reverse engineering the secret triggers embedded inside each client-submitted model at each round, unlearning them, and re-establishing the trigger-output associations in a harmless environment post-FL. However, referred to as the baseline method in this paper, it is inefficient as reverse engineering at each round incurs significant resource and time consumption, making it impractical for FL applications, as shown in Figure 3 (red). To optimize efficiency, we need to minimize server-side processing time in each round to prevent delay in subsequent rounds, and we also need to ensure a low resource consumption to enable parallel trigger recovery across multiple client-submitted models.
@@
Figure 3: Sanitizer offers significantly better scalability. It keeps the total server-side time consumption consistently lower (green) than the baseline (red) as the number of participating clients increases.
@@
In this paper, Sanitizer aims to efficiently sanitize backdoors for harmless client-side watermarking. To tackle the challenges mentioned above, we propose to first extract a small backdoor subnet from each client-submitted model for reverse engineering, driven by the observation that backdoors are often encapsulated in a subset of neurons. Then, we take advantage of the iterative nature in FL and propose a lightweight method to recover the trigger implanted in each client-submitted model by spreading the reverse optimization across communication rounds. The trigger and target class gradually emerge and take shape over time. Based on the above two improvements, Sanitizer is efficient, requires minimal GPU resources, and does not hamper the utility of the FL-trained model [24]. As shown in Figure 3, Sanitizer maintains consistently low time consumption (green). With only 25 clients, the baseline (red) takes approximately five times longer than Sanitizer. Given the fixed GPU memory of the server, the disparity between the two methods becomes more pronounced as the number of clients increases. Due to lower memory consumption, Sanitizer enables the server to process more client-submitted models in parallel. After completing the training, we employ a harmless relearning process to “unbind” the trigger from the clean images and re-establish its association with the unique, harmless environment of each client. Since each client-submitted model is processed independently before aggregation at each round and each client’s trigger is ultimately confined to its own client-specific harmless input subspace, watermark conflicts will not arise during the verification phase. In summary, our key contributions are as follows:
@@
  • We introduce a server-side sanitization process that mitigates potential harmful use in backdoor-based client-side watermarking by enabling the triggering of non-intuitive model behavior only in harmless environments.
@@ 2 Related Work
The increasing commercial and legal demands have fueled the development of watermarking techniques for protecting the IP of machine learning models [25, 26, 27, 8].
@@
Watermarking in Centralized Learning. Many watermarking approaches have been proposed to protect the IP of models trained via centralized learning. They can be categorized into two classes: (i) feature-based methods [28, 29, 30, 31, 32] that manipulate the loss function and model parameters and (ii) backdoor-based methods [33, 34, 35, 36, 37] that inject a trigger set with predefined output (known only to the model owner) as the watermark. These strategies for centralized learning lay the groundwork for the development of watermarking in FL, where the process is more complex due to the client-server learning protocol and specific privacy requirements [38, 39, 16].
@@
Watermarking in Federated Learning. Watermarking methods in FL can be categorized into server-side [39, 21] and client-side schemes, owing to the collaborative nature of client-server workflow [16]. We investigate the current client-side solutions of watermarking for IP protection in FL. One of the representative methods is FedIPR [11], which permits each client to embed their own backdoor-based and feature-based watermark into the model for contribution demonstration without exposing either their private watermark or training data to other clients. FedSOV [40] presents a cryptographic signature-based approach on top of FedIPR, allowing numerous clients to verify the ownership credentials through unforgeable digital signatures. FedCIP [38] is a feature-based watermarking framework that allows for traitor tracking while maintaining compatibility with secure aggregation [41]. Merkle-Sign [42] is also used for client-side watermarking, which designs a secure mechanism for distributed storage to protect both privacy and ownership. FedZKP [43] proposes a verification protocol based on the zero-knowledge proof for secure model ownership verification. Besides, [18] and [14] introduce a backdoor-based watermarking scheme for IP protection within a homomorphically encrypted FL framework using homomorphic encryption [41, 44] while preventing the watermarked updates from leakage.
@@
Research Gap. We argue that existing client-side approaches have certain limitations that warrant consideration to ensure comprehensive IP protection in FL. First, a similar backdoor trigger pattern with different target labels of any two clients is likely to cause collisions, resulting in both watermarks becoming unlearnable [17, 45]. Second, malicious clients can embed harmful misclassification rules as a malicious backdoor into the model under the guise of backdoor-based [20] watermarking, compromising the model’s integrity and posing a significant security threat [21, 46, 22, 47, 48] to real-world applications. To address these limitations, Sanitizer provides a client-agnostic server-side sanitization pipeline that not only makes the client-side watermarks harmless but also resolves potential conflicts.
@@ 3 Preliminaries
Backdoors can be utilized for both benign and malicious purposes. In the rest of this paper, when distinguishing between different purposes, we use the term “watermark” for backdoors serving benign purposes such as IP protection, and the term “harmful backdoor” or “malicious backdoor” for those serving malicious ones such as model control. Fundamentally, both purposes represent the same technique—a specific trigger-output pair embedded into a model. We use the term “backdoor” when uniformly referring to the backdoor itself without specifying purposes. These terms are used consistently throughout the paper.
In this section, we describe the threat model (Section 3.1), clarify the capabilities of Sanitizer (Section 3.2) and its goals (Section 3.3), and define the poisoning formulation and harmless setting in our context (Section 3.4).
@@ 3.1 Threat Model
We consider the threat model, where clients are authorized to embed legitimate backdoor-based watermarks into the model as proof of co-ownership or contribution during the local training phase. However, some clients are malicious. They exploit this privilege to unlawfully implant harmful backdoor rather than watermark, and finally, control the model predictions during the inference phase for harmful purposes. All clients follow the original FL protocol and do not cooperate with the server-side sanitization process.
@@ 3.2 Capabilities of Sanitizer
Consistent with prior works [39, 21, 49, 1], the server is considered a trusted and honest party employed by the collaborating clients to facilitate FL. Sanitizer runs on the server-side and has access to the compromised models received from clients. However, it cannot interfere with the standard local training process, modify client-side operations, or access the local dataset of individual clients. Moreover, Sanitizer has no prior knowledge about the potential backdoor triggers or target classes, and does not distinguish between benign and harmful backdoors. Similar to most existing defenses [50, 51, 52, 53, 54, 55], Sanitizer can only get access to a limited small portion of reserved clean dataset as defense data, which is common and only designed to drive the sanitization mechanisms. We exclude the defense data from all clients’ local training dataset, ensuring that there is no overlap between them.
@@ 3.3 Goals of Sanitizer
  • Effectiveness. Effectiveness implies that Sanitizer makes each client deterministically demonstrate their co-ownership or contribution to the FL-trained model by achieving near-perfect success in verifying the watermark in their respective harmless environment, while ensuring no conflicts arise among them.
@@
  • Fidelity. Fidelity refers to the requirement that Sanitizer should have negligible impact on the normal functionality of the FL-trained model, maintaining high accuracy on clean test samples.
  • Efficiency. The sanitization process should be efficient, with a low time and resource consumption on the server-side.
[…]
Figure 4: Examples of harmful (natural) inputs compared to the harmless (artificial) inputs of each client (e.g., Bob, Sam, and Jack).
@@ 3.4 Notations
in which $\mathcal{D}_{c}$ is the clean dataset of client $k$. $\mathcal{D}_{t}$ is the trigger-injected dataset for benign or malicious purposes. It is generated by placing a trigger pattern of its choice on clean data samples and flip their ground truths to be the malicious target label [20, 33, 56]. Besides, $f$ denotes the standard classifier parameterized by $\theta_{k}$. $\mathcal{L}$ denotes a standard loss function, e.g., cross-entropy.
Harmless Setting. We consider the harmless dataset $\mathcal{D}_{u}$ as a collection of unnatural images that lack practical significance and have no impact on the main task of the FL-trained model. Note that $\mathcal{X}_{u}$ must satisfy $\mathcal{X}_{u}\cap\mathcal{X}_{c}=\emptyset$, where $\mathcal{X}_{u}$ denotes the artificial input subspace and $\mathcal{X}_{c}$ denotes the natural input space, indicating no overlap between them, and ensuring $\mathcal{X}_{u}$ is distinct and significantly distant from the main task. In our setting, all the watermarks and backdoors should be transformed into a harmless environment (i.e., being treated equally), which is artificially and uniquely tailored for each client. For example, as shown in Figure 4, we use a set of distinct and unnatural inputs in unique colors as the harmless environments, compared to the harmful (natural) inputs.
@@ 4.1 Overview
Outline. Sanitizer is accomplished by three procedures during the FL process. First, we identify and extract the backdoor subnet (Section 4.2 and Section 4.3), followed by an efficient inversion of the backdoor trigger (Section 4.4), and simultaneously perform pruning and aggregation (Section 4.5) after the identification of backdoor subnet. After completing the training, we conduct harmless relearning (Section 4.6) to achieve the final harmless effectiveness of Sanitizer. We present the details in the following sections, exemplified using a specific client $k$ at a certain round $r+1$; the approach remains consistent and independent across all clients.